OpenStudy (anonymous):

what is the simplified form of the expression [m^-1 m^5/m^-2]^3?

OpenStudy (anonymous):

anything with a negative exponent in the numerator can be moved to the denominator. Likewise, a coefficient with a negative exponent in the denominator can be moved in to the numerator m^-1 = 1/m 1/m^-2 = m^2 Therefore, you have (m^5*m^2)/m = m^7/m = m^6 now plug this in to the original problem (m^6)^3 = m^18

OpenStudy (anonymous):

Okay give me a moment. @mramosville

OpenStudy (anonymous):

Actually I came up with the answer: 1/m^18 @mramosville

OpenStudy (triciaal):

when you multiply same base and the powers (m^-1)(m^5)(m^-2) =m^(-1+5-2) =m^2 when you raise a power to a power just multiply the powers so (m^2)^3 = m^(2)(3) = m^6

OpenStudy (triciaal):

correction the original problem had divide by m^-2 same principle would apply when you divide same base you subtract the powers

OpenStudy (triciaal):

(m^6)^3 = m^18